Should print VEDOS on the first two pages, light. Should print VEDOS at the bottom each page.

\usepackage[firsttwo,light,bottom]{draftcopy}

test6: This is a test page 1

This is a testpage 1 test6: This is a test page 2 This is a testpage 2 test6: This is a test page 3 This is a testpage 3 test6: This is a test page 4 This is a testpage 4